OpenStudy (brittneygarcia5):

Find the difference quotient and simplify: f(x)=x^2-x+1, (f(2+h)-f(2))/(h), h cannot equal 0

OpenStudy (brittneygarcia5):

\[f(x)=x ^{2}-x+1, \frac{ f(2+h)-f(2) }{ h }, h \neq0\]

satellite73 (satellite73):

slowly what is \(f(2)\)?

OpenStudy (brittneygarcia5):

is it \[f(x)\]

satellite73 (satellite73):

\[f(x)=x^2-x+1\] as you wrote so \[f(2)=2^2-2+1=?\]

OpenStudy (brittneygarcia5):

3

satellite73 (satellite73):

right good so next question, what is \(f(2+h)\)? hint, your answer will have an \(h\) in it hint hint: since \[f(\clubsuit)=\clubsuit^2-\clubsuit+1\] you have \[f(2+h)=(2+h)^2-(2+h)+1\] needs algebra to clean it up

OpenStudy (brittneygarcia5):

\[4+h^2-2-h+1=3+h^2-h\] Or would the h^2 and h turn into only one h?

satellite73 (satellite73):

close

satellite73 (satellite73):

\((2+h)\neq 4+h^2\) you got to multiply out \[(2+h)^2=(2+h)(2+h)=?\]

OpenStudy (brittneygarcia5):

4+4h+h^2?

satellite73 (satellite73):

yes

satellite73 (satellite73):

so now we are at \[4+4h+h^2-2-h+1-3\] for the top of the fraction \[f(2+h)-f(2)\] now just a tiny bit of arithemtetic

satellite73 (satellite73):

notice that the numbers add up to zero, all your are left with is an expression in \(H\)

satellite73 (satellite73):

\(h\)

satellite73 (satellite73):

oh and \(h^2\)

OpenStudy (brittneygarcia5):

So we're left with h^2+3h?

satellite73 (satellite73):

yes

satellite73 (satellite73):

that is the numerator so we are at \[\frac{3h+h^2}{h}\] as a fraction now divide each term by \(h\)

OpenStudy (brittneygarcia5):

So it's 3+h!

satellite73 (satellite73):

yes it is

satellite73 (satellite73):

this a calc class or not yet?

OpenStudy (brittneygarcia5):

Thank you so much, I struggled a lot with that problem. And it's for my pre-cal class

satellite73 (satellite73):

when you take calc by week 2 you till be able to come up the the answer instantly in your head

satellite73 (satellite73):

yw

OpenStudy (brittneygarcia5):

I hope so